OpenStudy (anonymous):

(1-3x/1+1/3x)/(1+3x/1-1/3x)

OpenStudy (pfenn1):

Is this right?\[\frac{1-\frac{3x}{1}+\frac{1}{3x}}{1+\frac{3x}{1}-\frac{1}{3x}}\]